OpenStudy (anonymous):

how to find slope perpendiclur to 4x-8y=8

OpenStudy (anonymous):

y=mx+c, then slope of the line perpendicular to this is , -1/m. so the slope of the perpendicular line to y=x/2+1 is, -2.

OpenStudy (anonymous):

??

OpenStudy (anonymous):

answer is -2. and the reason , i've given earlier, u got it?

OpenStudy (anonymous):

To find the slope of a line perpendicular to a given line, you take the inverse reciprocal. That's a fancy math way of saying you turn the fraction upside-down and change the sign. So, in your case you have the given line of \[4x-8y=8\]or, in slope-intercept form \[y=x/2-1\] Your slope is the coefficient of x, r, in this case, 1/2. Again, to find the slope of a perpendicular line, you take the inverse reciprocal (upside-down fraction with the sign changed). That would be -2.